# Unveiling the Role of Learning Rate Schedules via Functional Scaling Laws

### v1

@@ C.5.2 Upper and Lower Bounds
$\textbf{I}-\mathbf{H}^{\frac{1}{2}}\mathbf{W}^{\top}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})^{-1}\mathbf{W}\mathbf{H}^{\frac{1}{2}}$ $=\begin{pmatrix}\textbf{I}_{k}-\mathbf{H}_{0:k}^{1/2}\mathbf{W}_{0:k}^{\top}\mathbf{A}^{-1}\mathbf{W}_{0:k}\mathbf{H}_{0:k}^{1/2}-&-\mathbf{H}_{0:k}^{1/2}\mathbf{W}_{0:k}^{\top}\mathbf{A}^{-1}\mathbf{W}_{k:\infty}\mathbf{H}_{k:\infty}^{1/2}\\ -\mathbf{H}_{k:\infty}^{1/2}\mathbf{W}_{k:\infty}^{\top}\mathbf{A}^{-1}\mathbf{W}_{0:k}\mathbf{H}_{0:k}^{1/2}&\textbf{I}_{N-k}-\mathbf{H}_{k:\infty}^{1/2}\mathbf{W}_{k:\infty}^{\top}\mathbf{A}^{-1}\mathbf{W}_{k:\infty}\mathbf{H}_{k:\infty}^{1/2}\end{pmatrix}$
$=:\begin{pmatrix}\mathbf{X}&\mathbf{Y}\\ \mathbf{Y}^{\top}&\mathbf{Z}\end{pmatrix}$,
where $\mathbf{A}=\mathbf{W}\mathbf{H}\mathbf{W}^{\top}$.
We have
$\mathbb{E}_{\mathbf{u}_{0}}\|(\textbf{I}-\mathbf{H}^{\frac{1}{2}}\mathbf{W}^{\top}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})^{-1}\mathbf{W}\mathbf{H}^{\frac{1}{2}})\mathbf{H}^{\frac{1}{2}}\mathbf{u}_{0}\|^{2}$
$=\mathbb{E}_{\mathbf{u}_{0}}\left\langle\mathbf{X}^{2}+\mathbf{Y}\mathbf{Y}^{\top},\mathbf{H}_{0:k}^{1/2}\mathbf{u}_{0:k}\mathbf{u}_{0:k}^{\top}\mathbf{H}_{0:k}^{1/2}\right\rangle+\mathbb{E}_{\mathbf{u}_{0}}\left\langle\mathbf{Z}^{2}+\mathbf{Y}^{\top}\mathbf{Y},\mathbf{H}_{0:k}^{1/2}\mathbf{u}_{k:\infty}\mathbf{u}_{k:\infty}^{\top}\mathbf{H}_{k:\infty}^{1/2}\right\rangle$
$+2\mathbb{E}_{\mathbf{u}_{0}}\left\langle\mathbf{X}\mathbf{Y}+\mathbf{Y}\mathbf{Z},\mathbf{H}_{0:k}^{1/2}\mathbf{u}_{0:k}\mathbf{u}_{0:k}^{\top}\mathbf{H}_{k:\infty}^{1/2}\right\rangle$
$=\operatorname{Tr}((\mathbf{Z}^{2}+\mathbf{Y}\mathbf{Y}^{\top})\mathbf{H}_{0:k})+\operatorname{Tr}((\mathbf{Z}^{2}+\mathbf{Y}^{\top}\mathbf{Y})\mathbf{H}_{k:\infty}),$
We can calculate that
$\mathbf{Z}^{2}+\mathbf{Y}^{\top}\mathbf{Y}=\mathbf{Z}$
and
$\mathbf{X}^{2}+\mathbf{Y}\mathbf{Y}^{\top}=\mathbf{H}_{0:k}^{-1/2}\left[\mathbf{H}_{0:k}^{-1}+\mathbf{W}_{0:k}^{\top}\mathbf{A}_{k}^{-1}\mathbf{W}_{0:k}\right]^{-1}\mathbf{H}_{0:k}^{-1/2},$
where $\mathbf{A}_{k}=\mathbf{W}_{k:\infty}\mathbf{H}_{k:\infty}\mathbf{W}_{k:\infty}^{\top}$.
And then we have
$\operatorname{Tr}((\mathbf{X}^{2}+\mathbf{Y}\mathbf{Y}^{\top})\mathbf{H}_{0:k})+\operatorname{Tr}((\mathbf{Z}^{2}+\mathbf{Y}^{\top}\mathbf{Y})\mathbf{H}_{k:\infty})$
$=\operatorname{Tr}\left(\mathbf{H}_{0:k}^{-1/2}\left[\mathbf{H}_{0:k}^{-1}+\mathbf{W}_{0:k}^{\top}\mathbf{A}_{k}^{-1}\mathbf{W}_{0:k}\right]^{-1}\mathbf{H}_{0:k}^{-1/2}\mathbf{H}_{0:k}\right)+\operatorname{Tr}(\mathbf{Z}\mathbf{H}_{k:\infty})$
$=\operatorname{Tr}\left(\left[\mathbf{H}_{0:k}^{-1}+\mathbf{W}_{0:k}^{\top}\mathbf{A}_{k}^{-1}\mathbf{W}_{0:k}\right]^{-1}\right)+\operatorname{Tr}(\mathbf{Z}\mathbf{H}_{k:\infty})$
$\geqslant\operatorname{Tr}(\mathbf{Z}\mathbf{H}_{k:\infty}):=T_{3}$
We can then simplify $T_{3}$
$T_{3}$ $=\operatorname{Tr}\left(\mathbf{H}_{k:\infty}^{1/2}\left[\textbf{I}_{M-k}-\mathbf{H}_{k:\infty}^{1/2}\mathbf{W}_{k:\infty}^{\top}\mathbf{A}_{k}^{-1}\mathbf{W}_{k:\infty}\mathbf{H}_{k:\infty}^{1/2}\right]\right)$
$\geqslant\operatorname{Tr}\left(\mathbf{H}_{k:\infty}^{1/2}\left[\textbf{I}_{M-k}-\mathbf{H}_{k:\infty}^{1/2}\mathbf{W}_{k:\infty}^{\top}\mathbf{A}_{k}^{-1}\mathbf{W}_{k:\infty}\mathbf{H}_{k:\infty}^{1/2}\right]\mathbf{H}_{k:\infty}^{1/2}\right)$
where the second inequality follows from $A\succeq A_{k}$.
$T_{3}$ $\geqslant\sum_{i=1}^{N-k}\mu_{i}\left(\textbf{I}_{N-k}-\mathbf{H}_{k:\infty}^{1/2}\mathbf{W}_{k:\infty}^{\top}\mathbf{A}_{k}^{-1}\mathbf{W}_{k:\infty}\mathbf{H}_{k:\infty}^{1/2}\right)\cdot\mu_{N+1-k-i}(\mathbf{H}_{k:\infty})$
$\geqslant\sum_{i=1}^{N-k}\mu_{i}(\mathbf{M})\cdot\mu_{N+1-k-i}(\mathbf{H}_{k:\infty})\geqslant\sum_{i=k+M}^{N}\lambda_{i},$
where the second line follows from $A\succeq A_{k}$ and the third line follows from Von-Neuman’s Inequality. Since $\mathbf{M}=\mathbf{I}_{N-k}-\mathbf{H}_{k:\infty}^{1/2}\mathbf{W}_{k:\infty}^{\top}\mathbf{A}_{k}^{-1}\mathbf{W}_{k:\infty}\mathbf{H}_{k:\infty}^{1/2}$ is a projection matrix such that $\mathbf{M}^{2}=\mathbf{M}$ and $\operatorname{Tr}(\mathbf{I}_{N-k}-\mathbf{M})=N-k$, we have $\mathbf{M}$ has $M$ eigenvalues 0 and NkMN-k-M eigenvalues $1$.
Let $k$ be zero and substitute $\lambda_{i}\eqsim i^{-\beta}$ into the expression above, we have
$\sum_{i=M}^{N}\lambda_{i}\geqslant cM^{-s\beta}.$
@@
$\mathbf{A}^{\top}_{t}\mathbf{H}\mathbf{A}_{t}\succeq\mathbf{H}^{\frac{1}{2}}\mathbf{W}(\mathbf{I}-\mathbf{H}^{\frac{1}{2}}\mathbf{W}^{\top}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})^{-1}\mathbf{W}\mathbf{H}^{\frac{1}{2}})^{2}\mathbf{H}^{\frac{1}{2}}$
@@
First, we prove that
$\|\mathbf{M}_{t}\|_{2}\leqslant\frac{c}{t}.$
Note that the eigenvalues of $\mathbf{M}_{t}$ is $P_{t}(\hat{\lambda}_{j})=\frac{f(2t\hat{\lambda}_{j})}{2t}$, where
$f(x_{0})=x_{0}e^{-x_{0}}\leqslant\frac{1}{e}.$
So we have
$\|\mathbf{M}_{t}\|_{2}\leqslant\max_{1\leqslant j\leqslant M}P_{t}(\hat{\lambda}_{j})\leqslant\frac{1}{2et}.$
By definition of $T_{1}$, we have
@@
where the second line uses Woodbury’s identity. Since
@@
Assume $k\leqslant\frac{M}{2}$ and with probability at least $1-e^{-\Omega(M)}$ for some constant $c>0$ $\|\mathbf{W}_{0:k}\|_{2}\leqslant c$.
We may write $\mathbf{W}_{0:k}^{\top}\mathbf{A}_{k}^{-1}\mathbf{W}_{0:k}=\sum_{i=1}^{M}\frac{1}{\hat{\lambda}_{M-i}}\bm{s}_{i}\bm{s}_{i}^{\top}$, where $\bm{s}_{i}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,\mathbf{I}_{k}/N)$ and $(\hat{\lambda}_{i})_{i=1}^{M}$ are eigenvalues of $\mathbf{A}_{k}$ in non-increasing order. Therefore, for $k\leqslant M/3$,
$\sum_{i=1}^{M}\frac{1}{\hat{\lambda}_{M-i}}\bm{s}_{i}\bm{s}_{i}^{\top}\succeq\sum_{i=1}^{M/2}\frac{1}{\hat{\lambda}_{M-i}}\bm{s}_{i}\bm{s}_{i}^{\top}\succeq\frac{1}{\hat{\lambda}_{M/2}}\sum_{i=1}^{M/2}\bm{s}_{i}\bm{s}_{i}^{\top}\succeq\frac{c\mathbf{I}_{k}}{\hat{\lambda}_{M/2}}.$
@@
$\leqslant\frac{\mu_{N/2}(\mathbf{A}_{k})}{\mu_{N}(\mathbf{A}_{k})}.$
@@
$T_{2}$ $=\mathbf{u}_{k:\infty}^{*}{}^{\top}\mathbf{H}_{k:\infty}\mathbf{W}_{k:\infty}^{\top}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})^{-1/2}\exp(-2t\mathbf{W}\mathbf{H}\mathbf{W}^{\top})(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})^{-1/2}\mathbf{W}_{k:\infty}\mathbf{H}_{k:\infty}\mathbf{u}_{k:\infty}^{*}$
$\leqslant\mathbf{u}_{k:\infty}^{*}{}^{\top}\mathbf{H}_{k:\infty}\mathbf{W}_{k:\infty}^{\top}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})^{-1}\mathbf{W}_{k:\infty}\mathbf{H}_{k:\infty}\mathbf{u}_{k:\infty}^{*}$
$\leqslant\|\mathbf{H}_{k:\infty}^{1/2}\mathbf{W}_{k:\infty}^{\top}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})^{-1}\mathbf{W}_{k:\infty}\mathbf{H}_{k:\infty}^{1/2}\|\cdot\|\mathbf{u}_{k:\infty}^{*}\|^{2}_{\mathbf{H}_{k:\infty}}$
$\leqslant\|\mathbf{u}_{k:\infty}^{*}\|^{2}_{\mathbf{H}_{k:\infty}},$
@@
$\leqslant\|\mathbf{H}_{k:\infty}^{1/2}\mathbf{W}_{k:\infty}^{\top}\mathbf{A}_{k}^{-1}\mathbf{W}_{k:\infty}\mathbf{H}_{k:\infty}^{1/2}\|_{2}\leqslant 1.$
Combining Lemma C.17, C.18 and C.19, we get with probability at least $1-e^{-\Omega(M)}$,
$\mathbb{E}[\mathcal{E}_{t}]\eqsim g_{s}(t)+\int_{0}^{t}\mathcal{K}(t-\tau)\gamma_{\tau}(c_{\tau}\mathbb{E}[\mathcal{E}_{\tau}]+\sigma^{2})\mathop{}\!\mathrm{d}\tau+M^{-s\beta},$ (23)
which is of the same form as in Lemma C.12. From here following the same proof as before, we get the functional scaling laws for random projection matrices.
@@ D.1 Proofs for Constant LRS
When the learning rate $\eta(k)\equiv\eta$ and batch size $B$ are constants, for the top-$M$ selection of the projection matrix $\mathbf{W}$ or for the random case with probability at least $1-e^{-\Omega(M)}$, we have
@@
By our main Theorem 4.1, when the learning rate $\eta(k)\equiv\eta$, denote $\gamma:=\frac{\eta}{B}$, we have
$\mathbb{E}[\mathcal{E}_{K}]\eqsim(t)^{-s}+\gamma\int_{0}^{t}\mathcal{K}(t-r)(e(r)+\sigma^{2})\mathop{}\!\mathrm{d}\tau+M^{-s\beta}.$
Hence by Lemma C.9 and Lemma C.7, note that when $M^{\beta}<t$ we have $M^{-s\beta}\gtrsim\frac{1}{t^{s}}\gtrsim e(t)$, therefore
$\mathbb{E}[\mathcal{E}_{K}]\eqsim\frac{1}{t^{s}}+\frac{\gamma}{t^{s}}+\gamma\sigma^{2}+M^{-s\beta},$
@@ D.2 Proof for The Exponential-Decay LRS
$\varphi(\tau)=ae^{-\lambda\tau},\text{ with }\varphi(K)=b,$
where $\lambda=\log(a/b)/K=:1/\bar{K}$. Note that the intrinsic-time transform is given by
$T(\tau)=\int_{0}^{\tau}\varphi(r)\mathop{}\!\mathrm{d}r=\frac{a}{\lambda}\left(1-e^{-\lambda\tau}\right).$
@@
    $T(K)=\frac{a}{\lambda}(1-e^{-\lambda K})=\frac{K}{\log(a/b)}(a-b)=:\bar{K}(a-b).$
@@
$\gamma_{\varphi}(t)=\varphi(T^{-1}(t))=a-\lambda t.$
@@
The noise term satisfies $\mathcal{N}(\varphi)=bI_{1}+(a-b)I_{2}$ with
$I_{1}=\int_{M^{-\beta}}^{1}\frac{1-e^{-2uT}}{2u^{1/\beta}}\mathop{}\!\mathrm{d}u,\qquad I_{2}=\int_{M^{-\beta}}^{1}\left(\frac{1-e^{-2uT}-2uTe^{-2uT}}{4Tu^{1+1/\beta}}\right)\mathop{}\!\mathrm{d}u.$
@@
Noticing $b=a-\lambda T$ and $\lambda T=a-b$, we have
$\int_{0}^{T}\mathcal{K}(T-t)\gamma_{\varphi}(t)\mathop{}\!\mathrm{d}t$ $=\int_{0}^{T}\left(\int_{M^{-\beta}}^{1}u^{1-1/\beta}e^{-2u(T-t)}\mathop{}\!\mathrm{d}u\right)(a-\lambda t)\mathop{}\!\mathrm{d}t$
$=\int_{M^{-\beta}}^{1}u^{1-1/\beta}e^{-2uT}\left(\int_{0}^{T}e^{2ut}(a-\lambda t)\mathop{}\!\mathrm{d}t\right)\mathop{}\!\mathrm{d}u$
$=\int_{M^{-\beta}}^{1}u^{1-1/\beta}e^{-2uT}\left[\frac{a}{2u}\left(e^{2uT}-1\right)-\frac{\lambda}{2u}\left(Te^{2uT}-\frac{e^{2uT}-1}{2u}\right)\right]\mathop{}\!\mathrm{d}u$
$=\int_{M^{-\beta}}^{1}\left[\frac{a}{2u^{1/\beta}}-\frac{ae^{-2uT}}{2u^{1/\beta}}-\frac{\lambda T}{2u^{1/\beta}}+\frac{\lambda(1-e^{-2uT})}{4u^{1+1/\beta}}\right]\mathop{}\!\mathrm{d}u$
$=\int_{M^{-\beta}}^{1}\left[\frac{a-\lambda T}{2u^{1/\beta}}-\frac{(a-\lambda T+\lambda T)e^{-2uT}}{2u^{1/\beta}}+\frac{\lambda(1-e^{-2uT})}{4u^{1+1/\beta}}\right]\mathop{}\!\mathrm{d}u$
$=(a-\lambda T)\int_{M^{-\beta}}^{1}\frac{1-e^{-2uT}}{2u^{1/\beta}}\mathop{}\!\mathrm{d}u+\lambda T\int_{M^{-\beta}}^{1}\left(-\frac{e^{-2uT}}{2u^{1/\beta}}+\frac{1-e^{-2uT}}{4Tu^{1+1/\beta}}\right)\mathop{}\!\mathrm{d}u.$
@@
If $T$ and $M$ is sufficiently large, then $I_{1}=\frac{\beta}{2\beta-1}+o_{T,M}(1)$.
@@
Note that
$\int_{M^{-\beta}}^{1}\frac{1}{2u^{1/\beta}}\mathop{}\!\mathrm{d}u=\frac{\beta(1-M^{-(\beta-1)})}{2(\beta-1)}=:A.$
and
$\int_{M^{-\beta}}^{1}\frac{e^{-2uT}}{2u^{1/\beta}}\mathop{}\!\mathrm{d}u=\frac{1}{2(2T)^{1-1/\beta}}\int_{T/M^{\beta}}^{T}\frac{e^{-r}}{r^{1/\beta}}\mathop{}\!\mathrm{d}r\leqslant\frac{\Gamma(1+\frac{1}{\beta})}{2^{2-1/\beta}}\frac{1}{T^{1-1/\beta}}=:B.$
Then, we complete the proof by noting $I_{1}=A-B$. ∎
@@
If $T$ and $M$ is sufficiently large, then
$I_{2}\eqsim\frac{\beta\min(M,T^{1/\beta})}{4T}.$

### v3

@@ C.4.2 Upper and Lower Bounds
$T_{3}$ $=\left\langle\mathbf{I}_{N}-\mathbf{H}^{1/2}\mathbf{W}^{\top}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})^{-1}\mathbf{W}\mathbf{H}^{1/2},\,\mathbf{H}^{1/2}\mathbf{u}_{0}\mathbf{u}_{0}^{\top}\mathbf{H}^{1/2}\right\rangle,$
where the inner product $\langle\mathbf{A},\mathbf{B}\rangle=\operatorname{tr}(\mathbf{A}^{\top}\mathbf{B})$ denotes the trace inner product between matrices. Therefore note that $\mu_{i}(\mathbf{H}^{\frac{1}{2}}\mathbf{u}_{0}\mathbf{u}_{0}^{\top}\mathbf{H}^{\frac{1}{2}})=i^{-1-s\beta}$ by source and capacity conditions,
$T_{3}$ $\geqslant\sum_{i=1}^{N}\mu_{i}\left(\textbf{I}_{N}-\mathbf{H}^{1/2}\mathbf{W}^{\top}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})^{-1}\mathbf{W}\mathbf{H}^{1/2}\right)\cdot\mu_{N+1-i}(\mathbf{H}^{\frac{1}{2}}\mathbf{u}_{0}\mathbf{u}_{0}^{\top}\mathbf{H}^{\frac{1}{2}})$
$\gtrsim\sum_{i=1}^{N}\mu_{i}(\mathbf{M})\cdot(N+1-i)^{-1-s\beta}$
where the second line follows again from Von Neumann’s trace inequality. Since $\mathbf{M}=\mathbf{I}_{N}-\mathbf{H}^{1/2}\mathbf{W}^{\top}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})^{-1}\mathbf{W}\mathbf{H}^{1/2}$ is a projection matrix such that $\mathbf{M}^{2}=\mathbf{M}$ and $\operatorname{rank}(\mathbf{I}_{N}-\mathbf{M})=M$ with probability $1$, $\mathbf{M}$ must have $M$ eigenvalues 0 and NMN-M eigenvalues $1$.
Hence we have
$T_{3}\gtrsim\sum_{i=M}^{N}i^{-1-s\beta}\gtrsim M^{-s\beta}.$
@@
$\mathbf{A}^{\top}_{t}\mathbf{H}\mathbf{A}_{t}\succeq\mathbf{H}^{\frac{1}{2}}(\mathbf{I}-\mathbf{H}^{\frac{1}{2}}\mathbf{W}^{\top}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})^{-1}\mathbf{W}\mathbf{H}^{\frac{1}{2}})^{2}\mathbf{H}^{\frac{1}{2}}$
@@
Recall that the eigenvalues of $\mathbf{M}_{t}=P_{t}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})$ is $P_{t}(\widehat{\lambda}_{j})$, so we have
$\|\mathbf{M}_{t}\|_{2}=\max_{1\leqslant j\leqslant M}P_{t}(\widehat{\lambda}_{j})=\max_{1\leqslant j\leqslant M}\widehat{\lambda}_{j}e^{-2t\widehat{\lambda}_{j}}=\frac{1}{2t}\max_{1\leqslant j\leqslant M}2t\widehat{\lambda}_{j}e^{-2t\widehat{\lambda}_{j}}\leqslant\frac{1}{2t}\sup_{x>0}xe^{-x}\leqslant\frac{1}{2et},$
where the last step follows from $\sup_{x>0}xe^{-x}\leqslant e^{-1}$.
By the definition of $T_{1}$, we have
@@
where the first line uses Sherman–Morrison-Woodbury’s identity: For matrices $\mathbf{A}$, $\mathbf{C}$, $\mathbf{U}$, $\mathbf{V}$ with suitable shapes, we have
$(\mathbf{A}+\mathbf{U}\mathbf{C}\mathbf{V})^{-1}=\mathbf{A}^{-1}-\mathbf{A}^{-1}\mathbf{U}(\mathbf{C}^{-1}+\mathbf{V}\mathbf{A}^{-1}\mathbf{U})^{-1}\mathbf{V}\mathbf{A}^{-1}.$
Since
@@
Assume $k\leqslant\frac{M}{2}$ and with probability at least $1-e^{-\Omega(M)}$ for some constant $c>0$, $\|\mathbf{W}_{0:k}\|_{2}\leqslant c$.
We may write $\mathbf{W}_{0:k}^{\top}\mathbf{A}_{k}^{-1}\mathbf{W}_{0:k}=\sum_{i=1}^{M}\frac{1}{\widehat{\lambda}_{M-i}}\bm{s}_{i}\bm{s}_{i}^{\top}$, where $\bm{s}_{i}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,\mathbf{I}_{k}/N)$ and $(\widehat{\lambda}_{i})_{i=1}^{M}$ are eigenvalues of $\mathbf{A}_{k}$ in non-increasing order. Therefore, for $k\leqslant M/3$,
$\sum_{i=1}^{M}\frac{1}{\widehat{\lambda}_{M-i}}\bm{s}_{i}\bm{s}_{i}^{\top}\succeq\sum_{i=1}^{M/2}\frac{1}{\widehat{\lambda}_{M-i}}\bm{s}_{i}\bm{s}_{i}^{\top}\succeq\frac{1}{\widehat{\lambda}_{M/2}}\sum_{i=1}^{M/2}\bm{s}_{i}\bm{s}_{i}^{\top}\succeq\frac{c\mathbf{I}_{k}}{\widehat{\lambda}_{M/2}}.$
@@
$\leqslant\frac{\mu_{M/2}(\mathbf{A}_{k})}{\mu_{M}(\mathbf{A}_{k})}.$
@@
$T_{2}$ $=\mathbf{u}_{k:\infty}{}^{\top}\mathbf{H}_{k:\infty}\mathbf{W}_{k:\infty}^{\top}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})^{-1/2}\exp(-2t\mathbf{W}\mathbf{H}\mathbf{W}^{\top})(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})^{-1/2}\mathbf{W}_{k:\infty}\mathbf{H}_{k:\infty}\mathbf{u}_{k:\infty}$
$\leqslant\mathbf{u}_{k:\infty}{}^{\top}\mathbf{H}_{k:\infty}\mathbf{W}_{k:\infty}^{\top}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})^{-1}\mathbf{W}_{k:\infty}\mathbf{H}_{k:\infty}\mathbf{u}_{k:\infty}$
$\leqslant\|\mathbf{H}_{k:\infty}^{1/2}\mathbf{W}_{k:\infty}^{\top}(\mathbf{W}\mathbf{H}\mathbf{W}^{\top})^{-1}\mathbf{W}_{k:\infty}\mathbf{H}_{k:\infty}^{1/2}\|\cdot\|\mathbf{u}_{k:\infty}\|^{2}_{\mathbf{H}_{k:\infty}}$
$\leqslant\|\mathbf{u}_{k:\infty}\|^{2}_{\mathbf{H}_{k:\infty}},$
@@
$\leqslant\|\mathbf{H}_{k:\infty}^{1/2}\mathbf{W}_{k:\infty}^{\top}\mathbf{A}_{k}^{-1}\mathbf{W}_{k:\infty}\mathbf{H}_{k:\infty}^{1/2}\|_{2}=1.$
The last line is because a nonzero projection matrix has norm $1$. ∎
C.4.3 Proof of Proposition C.11
Combining Lemma C.14, C.15 and C.16, we get with probability at least $1-e^{-\Omega(M)}$,
$\mathbb{E}[\mathcal{E}_{t}]$ $\eqsim\mathbf{u}_{0}^{\top}\mathbf{A}_{t}^{\top}\mathbf{H}\mathbf{A}_{t}\mathbf{u}_{0}+\int_{0}^{t}\mathrm{tr}(\mathbf{S}\mathbf{A}_{t-\tau}^{\top}\mathbf{H}\mathbf{A}_{t-\tau}\mathbf{S})\gamma_{\tau}(\mathbb{E}[\mathcal{E}_{\tau}]+\sigma^{2})\mathop{}\!\mathrm{d}\tau,$ (43)
$\eqsim M^{-s\beta}+\widehat{e}_{M}(t)+\int_{0}^{t}\widehat{\mathcal{K}}_{M}(t-\tau)\gamma_{\tau}(\mathbb{E}[\mathcal{E}_{\tau}]+\sigma^{2})\mathop{}\!\mathrm{d}\tau,$ (44)
Here we arrive at the Volterra equation for random feature case. Since we have $\widehat{\lambda}_{j}\eqsim\lambda_{j}$, the proof in the top-$M$ case also holds for $\widehat{e}_{M}$ and $\widehat{\mathcal{K}}_{M}$.
@@ Appendix D Proofs for Section 5
When the condition $\sigma\gtrsim 1$ holds – indicating a constant label-noise level – the FSL simplifies to
$\mathbb{E}[\mathcal{E}(\bm{\nu}_{t})]\eqsim\frac{1}{M^{s\beta}}+\frac{1}{t^{s}}+\frac{\sigma^{2}}{B}\mathcal{N}(\varphi),\quad\text{ with }\quad\mathcal{N}(\varphi)=\int_{0}^{t}\mathcal{K}_{M}(t-r)\varphi(T^{-1}(r))\mathop{}\!\mathrm{d}r,$ (45)
where the fit-dependent noise term $e(r)$ is absorbed by the full-batch GD term due to $\sigma\gtrsim 1$. Extending to the full range $\sigma\geqslant 0$ is possible but makes the statements and derivations much more involved. We therefore focus on the above case to streamline the exposition.
@@ D.1 Proofs for Constant LRS
Under Assumption 5.1, when the learning rate $\eta(k)\equiv\eta$, for the top-$M$ selection of the projection matrix $\mathbf{W}$ or for the random case with probability at least $1-e^{-\Omega(M)}$, we have
@@
By our main Theorem 4.2, when the learning rate $\eta(k)\equiv\eta$, denote $\gamma:=\frac{\eta}{B}$, by (45) we have
@@ D.2 Proof for The Exponential-Decay LRS
$\varphi(\tau)=ae^{-\zeta\tau},\text{ with }\varphi(K)=b,$
where $\zeta=\log(a/b)/K=:1/\bar{K}$. Note that the intrinsic-time transform is given by
$T(\tau)=\int_{0}^{\tau}\varphi(r)\mathop{}\!\mathrm{d}r=\frac{a}{\zeta}\left(1-e^{-\zeta\tau}\right).$
@@
    $T(K)=\frac{a}{\zeta}(1-e^{-\zeta K})=\frac{K}{\log(a/b)}(a-b)=:\bar{K}(a-b).$
@@
$\gamma(t)=\varphi(T^{-1}(t))=a-\zeta t.$
@@
When $T\gtrsim 1$, the noise term satisfies
$\mathcal{N}(\varphi):=\int_{0}^{T}\mathcal{K}_{M}(T-t)\gamma(t)\mathop{}\!\mathrm{d}t=bI_{1}+(a-b)I_{2},$
where
$I_{1}=\sum_{j=1}^{M}\frac{1-e^{-2\lambda_{j}T}}{2}\lambda_{j},\qquad I_{2}=\sum_{j=1}^{M}\left(-\frac{\lambda_{j}e^{-2\lambda_{j}T}}{2}+\frac{1-e^{-2\lambda_{j}T}}{4T}\right).$
@@
We approximate the forgetting kernel by summation:
$\mathcal{K}_{M}(t)=\sum_{j=1}^{M}\lambda_{j}^{2}e^{-2\lambda_{j}t}.$
Noticing $b=a-\zeta T$ and $\zeta T=a-b$, we have
$\int_{0}^{T}\mathcal{K}_{M}(T-t)\gamma(t)\mathop{}\!\mathrm{d}t$ $=\sum_{j=1}^{M}\lambda_{j}^{2}\int_{0}^{T}e^{-2\lambda_{j}(T-t)}(a-\zeta t)\mathop{}\!\mathrm{d}t$
$=\sum_{j=1}^{M}\lambda_{j}^{2}e^{-2\lambda_{j}T}\left[a\!\int_{0}^{T}e^{2\lambda_{j}t}\mathop{}\!\mathrm{d}t-\zeta\!\int_{0}^{T}te^{2\lambda_{j}t}\mathop{}\!\mathrm{d}t\right]$
$=\sum_{j=1}^{M}\left[\frac{a(1-e^{-2\lambda_{j}T})}{2}\lambda_{j}-\frac{\zeta(1-(1+2\lambda_{j}T)e^{-2\lambda_{j}T})}{4}\right]$
$=(a-\zeta T)\sum_{j=1}^{M}\frac{1-e^{-2\lambda_{j}T}}{2}\lambda_{j}+\zeta T\sum_{j=1}^{M}\left(-\frac{\lambda_{j}e^{-2\lambda_{j}T}}{2}+\frac{1-e^{-2\lambda_{j}T}}{4T}\right).$
@@
If $T$ and $M$ are sufficiently large, then
$I_{1}=C_{\beta}+o_{T,M}(1),$
where $C_{\beta}$ is a constant depending only on $\beta$.
@@
Recall that
$I_{1}=\sum_{j=1}^{M}\frac{1-e^{-2\lambda_{j}T}}{2}\lambda_{j}=\frac{1}{2}\sum_{j=1}^{M}\lambda_{j}-\frac{1}{2}\sum_{j=1}^{M}\lambda_{j}e^{-2\lambda_{j}T}=:A-B.$
For the first term,
$A=\frac{1}{2}\sum_{j=1}^{M}\lambda_{j}=C_{\beta}-\frac{1}{2}\sum_{j=M+1}^{\infty}\lambda_{j},\qquad C_{\beta}:=\frac{1}{2}\sum_{j=1}^{\infty}\lambda_{j}.$
By the integral bound for the tail,
$\sum_{j=M+1}^{\infty}\lambda_{j}\leqslant c\sum_{j=M+1}^{\infty}j^{-\beta}\leqslant c\int_{M}^{\infty}x^{-\beta}\,dx=\frac{cM^{-(\beta-1)}}{\beta-1},$
so $A=C_{\beta}+O(M^{-(\beta-1)})=C_{\beta}+o_{M}(1)$.
For the second term, following the same proof in Lemma C.4, we have
$B=\frac{1}{2}\sum_{j=1}^{M}\lambda_{j}e^{-2\lambda_{j}T}\lesssim T^{-(1-1/\beta)}.$
Combining the two estimates gives
$I_{1}=C_{\beta}+o_{M}(1)+O(T^{-(1-1/\beta)})=C_{\beta}+o_{T,M}(1).$
@@
If $T$ and $M$ are sufficiently large, then
$I_{2}\eqsim\frac{\min(M,\,T^{1/\beta})}{T}.$